OpenStudy (anonymous):

The normal line to a curve C at a point P is, by definition, the line that passes through P and is perpendicular to the tangent line to C at P. Find the equation of the normal line to the parabola y=−3−x^2 at the point (2,−7) .

OpenStudy (anonymous):

can you find what is the slope of the parabola at the given point?

OpenStudy (anonymous):

I don't know. I don't know how to even approach this problem!

OpenStudy (anonymous):

what does the slope of a line represent in general terms?

OpenStudy (anonymous):

its the rate of change, do you know any other function that represents rate or change?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

derivatives mean rate of change

OpenStudy (anonymous):

does that mean I need to find the derivative of the given function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and that's -2x

OpenStudy (anonymous):

so thats the slope of the parabola, can you find the equation of the tangent at the point?

OpenStudy (anonymous):

is this where I use y-y=m(x-x)

OpenStudy (anonymous):

yeah, that is the point slope formula.

OpenStudy (anonymous):

is this right? y-(-7)=-2x(x-2)

OpenStudy (anonymous):

no, the slope in general is -2x, but you need the slope at a certain point.

OpenStudy (anonymous):

Can you just tell me step by step on how to do this?

OpenStudy (anonymous):

so the slope at any point on the par is -2x but you know what x you are interested in right? x=2 so the slope there is -4

OpenStudy (anonymous):

so now do I just plug the -4 into the slope formula?

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

y-(-7)=-4(x-2)?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

sorry, i'm not getting notifications. The slope of the tangent at that point is -4, you wan the equation of the normal line. like the question says the normal line is perpendicular. that means it has an inverse slope. do you know what the inverse of -4 is?

OpenStudy (anonymous):

the answer to this question is due at 12 tonight. Can you just give me the answer and the after we can talk about why its the answer?

OpenStudy (anonymous):

its only a few steps

OpenStudy (anonymous):

then tell me the steps pleaseeeeeeeeeee

OpenStudy (anonymous):

you need the inverse of the slope to get the perpendicular slope and you already know the point so you can use the point slope formula

OpenStudy (anonymous):

whats the inverse of the slope?

OpenStudy (anonymous):

 does that help?